# Primes and almost primes between cubes

### v1

@@ Abstract.
The first author was supported by Australian Research Council Discovery Project DP240100186 and an Australian Mathematical Society Lift-off Fellowship. The second author was supported by an Australian Government Research Training Program (RTP) Scholarship. Computing Infrastructure at Butler University was supported by Frank Levinson and the Fairbanks Foundation. Computing Facilities were also provided by the National Computational Infrastructure of Australia.
@@ 1.1  Overview
As a first step towards resolving Conjecture 1.2 unconditionally, we consider how far one can go with modern computations. Here, state-of-the-art interval estimates [7, Table 2] imply that there is always a prime between $n^{3}$ and $(n+1)^{3}$ for $n^{3}\leq 1.42\cdot 10^{35}$. However, by adapting a recent algorithm of the third and fourth authors [21], we are able to improve this result by a factor of $10^{5}$.
@@
To prove Theorem 1.4, we utilise a logarithmic sieve weighting procedure attributed to Richert [19]. By contrast, in [9] a simpler weighting procedure due to Kuhn [16] was used. We found, more than we expected, that Richert’s logarithmic weights give much better numerical results than Kuhn’s weighting procedure. For instance, we performed some rough calculations using Kuhn’s weighting procedure and were not able to prove Theorem 1.4, even for $n\geq 10^{60}$. We also note that other more complicated weighting procedures exist that improve upon Richert’s weights asymptotically (see e.g. [12, §5.2–5.3]). However, given the complexity of these weighting procedures, it is not clear whether they would yield better numerical results than Richert’s weights, except for when dealing with very large numbers.
@@
It is conjectured that $f(n)$ is prime for all $n\geq 1$ and this would essentially follow from Conjecture 1.2. We do not give the proof of Corollary 1.5 here, since the argument is identical to that in Mill’s original paper [17], combined with the computations in [5].
Finally, we remark that using our approach it is not possible to improve Theorem 1.4 to $\Omega(a)=1$, even for sufficiently large $n$. Rather, to detect primes between consecutive cubes using known sieve methods, one would require a more powerful tool. Namely, one would need either an explicit version of Harman’s prime-detecting sieve (see [13, Chapter 7]) or an explicit version of Iwaniec’s linear sieve with a well-factorable remainder term (see [12, Chapter 6]), neither of which have been established in the literature.
@@ 3.1  An explicit linear sieve
$h(s):=\begin{cases}s^{-1}e^{-2},&0\leq s\leq 1,\\ e^{-2}&1\leq s\leq 2,\\ e^{-s}&2\leq s\leq 3,\\ 3s^{-1}e^{-s}&s\geq 3,\end{cases}$ (3.7)
@@
From (3.8), we have the following expressions for $F(s)$ when $0\leq s\leq 3$ and $f(s)$ when $2\leq s\leq 4$, which are sufficient ranges of $s$ to consider for our application.
@@
We remark that the definition for $h(s)$ when $0\leq s\leq 1$ in (3.7) is nonstandard and in fact does not appear in the statement of [3, Theorem 2.1]. This is because this range of $s$ is outside the limits of where the sieve upper and lower bounds, (3.5) and (3.6), can be applied. However, setting $h(s)=s^{-1}e^{-2}$ for $0\leq s\leq 1$ is convenient notation for our application, since we will require a (weak) upper bound for $S(\operatorname{\mathcal{A}},\operatorname{\mathcal{P}},z)$ that essentially extends (3.5) to the case $D<z$.
The density function $g(d)$ in Lemma 3.1 should be chosen so that $|\operatorname{\mathcal{A}}|\approx g(d)|\operatorname{\mathcal{A}}_{d}|$, yielding a small value for $|r(d)|$. Since for our application, $\operatorname{\mathcal{A}}$ is just a set of consecutive integers, we choose $g(d)=\frac{1}{d}$ so that
@@
Before continuing, we also give two explicit versions of (3.4) with $g(p)=1/p$, which will suffice for our application.
@@
For all $u\geq 3$ and $z\geq 3024$, we have
$\prod_{u\leq p<z}\left(1-\frac{1}{p}\right)^{-1}<(1+1.97\cdot 10^{-3})\frac{\log z}{\log u}.$ (3.9)
Then, for all $u\geq 3$ and $z\geq 10^{5}$, we have
@@
We begin by noting that (3.9) is [9, Proposition A.2] and the proof of (3.10) is very similar. We thus omit the proof of (3.9) and outline the proof of (3.10), referring the reader to the appendix of [9] for more details.
For the proof of (3.10), we consider three cases for $z$.
@@
This case is identical to Case 3 in the proof of [9, Proposition A.2], whereby (3.10) follows from (3.11) and further bounds on $\prod_{p\leq x}\left(1-1/p\right)^{-1}$ due to Vanlalngaia [22, Theorem 5]. ∎
In the notation of Lemma 3.1, the inequality (3.9) says that one may take $Q=2$ and $\varepsilon=1.97\cdot 10^{-3}$} when $z\geq 3024$, and (3.10) says that one may take $Q=2$ and $\varepsilon=2.8\cdot 10^{-4}$ when $z\geq 10^{5}$.
@@ 3.2  Richert’s weighted sieve
$\sum_{\begin{subarray}{c}z\leq p<y\\ p\in\mathcal{P}\end{subarray}}|A_{p^{2}}|$
@@
Let $a^{\prime}\in\operatorname{\mathcal{A}}^{\prime}$ with $(a,P(z))=1$. For any prime factor $p$ of $a^{\prime}$, if $p^{2}|a^{\prime}$, then one must have $p\geq y$ and therefore $1-\frac{\log p}{\log y}\leq 0$. In particular,
@@
Our next goal is to relate $W(\operatorname{\mathcal{A}}^{\prime},\operatorname{\mathcal{P}},z)$ to $S(A,\operatorname{\mathcal{P}},z)$, so that $r_{k}(\operatorname{\mathcal{A}})$ may be bounded using Lemmas 3.1 and 3.5. In what follows, we will assume the condition
@@
$=\frac{1}{k}\left(\sum_{\begin{subarray}{c}a\in\operatorname{\mathcal{A}}\\ (a,P(z))=1\end{subarray}}w(a)-\sum_{\begin{subarray}{c}z\leq p<y\\ p\in\operatorname{\mathcal{P}}\end{subarray}}\sum_{\begin{subarray}{c}a\in\operatorname{\mathcal{A}}_{p^{2}}\\ (a,P(z))=1\end{subarray}}w(a)\right).$ (3.23)
@@ 3.3  An upper bound for the sum over $S(\operatorname{\mathcal{A}}_{p},\operatorname{\mathcal{P}},z)$
$w:=X^{\frac{2}{3}-\alpha-\frac{1}{k_{1}}}.$
@@
noting that $S(\operatorname{\mathcal{A}},\operatorname{\mathcal{P}},z)$ sifts out more elements of $\operatorname{\mathcal{A}}$ than $S(\operatorname{\mathcal{A}},\operatorname{\mathcal{P}},D_{q})$.
@@
$\leq k_{1}e^{-\gamma}\left(1+\frac{1}{2(\log\widetilde{D})^{2}}\right)\sum_{\begin{subarray}{c}q\in\operatorname{\mathcal{P}}\\ z\leq q\leq y\end{subarray}}\left(1-\frac{\log q}{\log y}\right)|\operatorname{\mathcal{A}}_{q}|\left(\frac{F(s_{q})+\varepsilon C_{1}(\varepsilon)e^{2}h(s_{q})}{{\log X}}\right)$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\sum_{\begin{subarray}{c}q\in\operatorname{\mathcal{P}}\\ z\leq q\leq w\end{subarray}}\left(1-\frac{\log q}{\log y}\right)\sum_{\begin{subarray}{c}d\mid P(z)\\ d<QD_{q}\end{subarray}}|r_{q}(d)|.$ (3.38)
@@
since $\operatorname{\mathcal{A}}$ is a set of consecutive integers. Now, because $k_{1}\leq 6$, we have $0\leq s_{q}\leq 3$ and thus by Lemma 3.2,
@@
$\qquad\leq\frac{1}{9(2+3\alpha)\log X}\left(3\log\frac{k_{1}}{k_{2}}+(k_{2}(2+3\alpha)-3)\log\frac{2+3\alpha-\frac{3}{k_{2}}}{2+3\alpha-\frac{3}{k_{1}}}\right)$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\frac{5k_{1}^{4}}{k_{\alpha}(\log X)^{3}}\left(1-\frac{k_{2}}{k_{1}}\right).$
@@
$\sum_{\begin{subarray}{c}q\in\operatorname{\mathcal{P}}\\ z\leq q<y\end{subarray}}\left(1-\frac{\log q}{\log y}\right)|\operatorname{\mathcal{A}}_{q}|\left(\frac{F(s_{q})+\varepsilon C_{1}(\varepsilon)e^{2}h(s_{q})}{\log X}\right)\leq M_{1}(X),$
@@
In this section we combine the results of Section 3 to prove Theorem 1.4. Here, the overarching approach is to use the computation in Theorem 1.3 for small $n$, and Lemma 3.6 for large $n$. From here onwards, the set $\operatorname{\mathcal{A}}$ is as defined in (3.1), $X$ is as in (3.13), and in the notation of Section 3, we set $k=2$, $k_{1}=5$, and $k_{2}=1.85$.
@@
For $a\geq 12$, the lemma follows directly from [11, Lemma 7]. The case $0<a\leq 12$} can then be verified via a short computation, noting that
@@
Our explicit version of ($Q_{0}(c,\delta)$) is then as follows. In particular, for $N>10^{40}$ we may take $c=0.03$ and $\delta=0.18$. This result could certainly be optimised further, however any improvement would have a negligible impact in our subsequent proof of Theorem 1.4.
@@
Let $N\geq 10^{40}$, $z=X^{1/5}$ and $y=X^{1/1.85}$. Then,
$\sum_{\begin{subarray}{c}z\leq q<y\\ q\in\mathcal{P}\end{subarray}}|\operatorname{\mathcal{A}}_{q^{2}}|\leq 0.03|\operatorname{\mathcal{A}}|^{0.82}.$
@@
$z\geq N^{1/5}=\frac{1}{3^{3/10}}(3N^{2/3})^{3/10}\geq\frac{1}{3^{3/10}}|\operatorname{\mathcal{A}}|^{3/10}.$
@@
$\frac{2.22|\operatorname{\mathcal{A}}|}{z\log z}\leq 2.22\cdot 3^{3/10}\frac{|\operatorname{\mathcal{A}}|^{7/10}}{\log(10^{8})}\leq 0.17|\operatorname{\mathcal{A}}|^{7/10}.$ (4.2)
@@
$\pi(y)\leq 1.04\frac{y}{\log y}\leq 0.021y\leq 0.021|\operatorname{\mathcal{A}}|^{0.82}.$ (4.3)
@@
$\sum_{\begin{subarray}{c}z\leq q<y\\ q\in\mathcal{P}\end{subarray}}|\operatorname{\mathcal{A}}_{q^{2}}|\leq 0.17|\operatorname{\mathcal{A}}|^{7/10}+0.021|\operatorname{\mathcal{A}}|^{0.82}\leq 0.03|\operatorname{\mathcal{A}}|^{0.82}$
@@
Let $N=n^{3}\geq 10^{40}$ throughout since smaller values of $N$ are covered by Theorem 1.3. We apply Lemma 4.3 and Lemma 3.6 with $k=2$, $z=X^{1/5}$, $y=X^{1/1.85}$}, and
$\lambda=k+1-k_{2}=1.15$
@@
$r_{2}(\operatorname{\mathcal{A}})\geq\frac{1.15}{2}S(\operatorname{\mathcal{A}},\operatorname{\mathcal{P}},z)-\frac{1}{2}\sum_{z\leq q\leq y}\left(1-\frac{\log q}{\log y}\right)S(\operatorname{\mathcal{A}}_{q},\operatorname{\mathcal{P}},z)-\frac{0.0345|\operatorname{\mathcal{A}}|^{0.82}}{2}.$ (4.4)
@@
$\frac{0.0345|\operatorname{\mathcal{A}}|^{0.82}}{2}\leq 10^{-4}\thinspace\frac{N^{2/3}}{\log X}.$ (4.5)
Next we give a lower bound for the first term in (4.4). Since $z=X^{1/5}>10^{5}$, applying Lemma 3.1 with Lemmas 3.3 and 3.4 gives
$S(\operatorname{\mathcal{A}},\operatorname{\mathcal{P}},z)$ $>|\operatorname{\mathcal{A}}|V(z)\cdot(f(s)-110\thinspace\varepsilon C_{2}(\varepsilon)e^{2}h(s))-\sum_{d<2D}\mu^{2}(d).$
@@
$f(s)-110\varepsilon e^{2}h(s)>C(s):=\frac{2e^{\gamma}\log(s-1)}{s}-0.0308e^{2-s}.$ (4.7)
To bound the remainder term in (4), we apply the bound (see [18, Lemma 4.6])
@@
noting that $D=z^{s}\geq z^{2}\geq 10^{16}$. Substituting (4.7) and (4.8) into (4),
@@
Since $k_{1}=5$, we find that the lower bound in (4.9) is maximised around $s=2.98$ whereby $C(s)>0.8$ and
$\frac{1.15}{2}S(\operatorname{\mathcal{A}},\operatorname{\mathcal{P}},z)>3.77\frac{N^{2/3}}{\log X}.$ (4.10)
Finally we bound the second term in (4.4). In the notation of Proposition 3.9 we set $\alpha=0.03$ so that
$D_{y}=X^{\frac{2}{3}-0.03-\frac{1}{1.85}}>6999$ (4.11)
and it is valid, via Lemma 3.4, to set $Q=2$ and $\varepsilon=1.97\cdot 10^{-3}$. In particular, Proposition 3.9 gives
$\frac{1}{2}\sum_{\begin{subarray}{c}q\in\operatorname{\mathcal{P}}\\ z\leq q<y\end{subarray}}\left(1-\frac{\log q}{\log y}\right)S(\operatorname{\mathcal{A}}_{q},\operatorname{\mathcal{P}},z)\leq\frac{5e^{-\gamma}}{2}\left(1+\frac{1}{2(\log D_{y})^{2}}\right)(M_{1}(X)+M_{2}(X))+\frac{\mathcal{E}(X)}{2}.$
@@
$\frac{5e^{-\gamma}}{2}\left(1+\frac{1}{(\log D_{y})^{2}}\right)\leq 1.43,$
$M_{1}(X)\leq\frac{0.78N^{2/3}}{\log X},$
$M_{2}(X)\leq(4\cdot 10^{-5})\frac{N^{2/3}}{\log X}$
@@
$r_{2}(\operatorname{\mathcal{A}})>0.53\frac{N^{2/3}}{\log X},$

### v2

@@ Abstract.
The first author was supported by Australian Research Council Discovery Project DP240100186 and an Australian Mathematical Society Lift-off Fellowship. The third author was supported by an Australian Government Research Training Program (RTP) Scholarship. Computing Infrastructure at Butler University was supported by Frank Levinson and the Fairbanks Foundation. Computing Facilities were also provided by the National Computational Infrastructure of Australia.
@@ 1.1  Overview
As a first step towards resolving Conjecture 1.2 unconditionally, we consider how far one can go with modern computations. Here, state-of-the-art interval estimates [7, Table 2] imply that there is always a prime between $n^{3}$ and $(n+1)^{3}$ for $n^{3}\leq 1.42\cdot 10^{35}$. However, by adapting a recent algorithm of the second and fourth authors [22], we are able to improve this result by a factor of $10^{5}$.
@@
To prove Theorem 1.4, we utilise a logarithmic sieve weighting procedure attributed to Richert [20]. By contrast, in [9] a simpler weighting procedure due to Kuhn [17] was used. We found, more than we expected, that Richert’s logarithmic weights give much better numerical results than Kuhn’s weighting procedure. For instance, we performed some rough calculations using Kuhn’s weighting procedure and were not able to prove Theorem 1.4, even for $n\geq 10^{60}$. We also note that other more complicated weighting procedures exist that improve upon Richert’s weights asymptotically (see e.g. [13, §5.2–5.3]). However, given the complexity of these weighting procedures, it is not clear whether they would yield better numerical results than Richert’s weights, except for when dealing with very large numbers.
@@
It is conjectured that $f(n)$ is prime for all $n\geq 1$ and this would essentially follow from Conjecture 1.2. We refer the reader to [11] for further discussion of unconditional variants of Mill’s constant. Note that we do not give the proof of Corollary 1.5 here, since the argument is identical to that in Mill’s original paper [18], combined with the computations in [5].
Finally, we remark that using our approach it is not possible to improve Theorem 1.4 to $\Omega(a)=1$, even for sufficiently large $n$. Rather, to detect primes between consecutive cubes using known sieve methods, one would require a more powerful tool. Namely, one would need either an explicit version of Harman’s prime-detecting sieve (see [14, Chapter 7]) or an explicit version of Iwaniec’s linear sieve with a well-factorable remainder term (see [13, Chapter 6]), neither of which have been established in the literature.
@@ 3.1  An explicit linear sieve
$h(s):=\begin{cases}s^{-1}e^{-2},&0<s\leq 1,\\ e^{-2}&1\leq s\leq 2,\\ e^{-s}&2\leq s\leq 3,\\ 3s^{-1}e^{-s}&s\geq 3,\end{cases}$ (3.7)
@@
From (3.8), we have the following expressions for $F(s)$ when $0<s\leq 3$ and $f(s)$ when $2\leq s\leq 4$, which are sufficient ranges of $s$ to consider for our application.
@@
We remark that the definition for $h(s)$ when $0<s\leq 1$ in (3.7) is nonstandard and in fact does not appear in the statement of [3, Theorem 2.1]. This is because this range of $s$ is outside the limits of where the sieve upper and lower bounds, (3.5) and (3.6), can be applied. However, setting $h(s)=s^{-1}e^{-2}$ for $0<s\leq 1$ is convenient notation for our application, since we will require a (weak) upper bound for $S(\operatorname{\mathcal{A}},\operatorname{\mathcal{P}},z)$ that essentially extends (3.5) to the case $D<z$.
The density function $g(d)$ in Lemma 3.1 should be chosen so that $|\operatorname{\mathcal{A}}_{d}|\approx g(d)|\operatorname{\mathcal{A}}|$, yielding a small value for $|r(d)|$. Since for our application, $\operatorname{\mathcal{A}}$ is just a set of consecutive integers, we choose $g(d)=\frac{1}{d}$ which gives
@@
Before continuing, we give an explicit version of (3.4) with $g(p)=1/p$ and a range of $z$ that suffices for our application.
@@
For all $u\geq 3$ and $z\geq 10^{5}$, we have
@@
The proof of (3.9) is very similar to [9, Proposition A.2]. We thereby omit some details for brevity, referring the reader to the appendix of [9].
For the proof, we consider three cases for $z$.
@@
This case is identical to Case 3 in the proof of [9, Proposition A.2], whereby (3.9) follows from (3.10) and further bounds on $\prod_{p\leq x}\left(1-1/p\right)^{-1}$ due to Vanlalngaia [23, Theorem 5]. ∎
In the notation of Lemma 3.1, the inequality (3.9) says that one may take $Q=2$ and $\varepsilon=2.8\cdot 10^{-4}$ when $z\geq 10^{5}$.
@@ 3.2  Richert’s weighted sieve
$\sum_{\begin{subarray}{c}z\leq p<y\\ p\in\mathcal{P}\end{subarray}}|\operatorname{\mathcal{A}}_{p^{2}}|$
@@
Let $a^{\prime}\in\operatorname{\mathcal{A}}^{\prime}$ with $(a^{\prime},P(z))=1$. For any prime factor $p$ of $a^{\prime}$, if $p^{2}|a^{\prime}$, then one must have $p\geq y$ and therefore $1-\frac{\log p}{\log y}\leq 0$. In particular,
@@
Our next goal is to relate $W(\operatorname{\mathcal{A}}^{\prime},\operatorname{\mathcal{P}},z)$ to $S(\operatorname{\mathcal{A}},\operatorname{\mathcal{P}},z)$, so that $r_{k}(\operatorname{\mathcal{A}})$ may be bounded using Lemmas 3.1 and 3.5. In what follows, we will assume the condition
@@
$\geq\frac{1}{k}\left(\sum_{\begin{subarray}{c}a\in\operatorname{\mathcal{A}}\\ (a,P(z))=1\end{subarray}}w(a)-\sum_{\begin{subarray}{c}z\leq p<y\\ p\in\operatorname{\mathcal{P}}\end{subarray}}\sum_{\begin{subarray}{c}a\in\operatorname{\mathcal{A}}_{p^{2}}\\ (a,P(z))=1\end{subarray}}w(a)\right).$ (3.22)
@@ 3.3  An upper bound for the sum over $S(\operatorname{\mathcal{A}}_{p},\operatorname{\mathcal{P}},z)$
$w:=X^{\frac{2}{3}-\alpha-\frac{1}{k_{1}}}$
@@
noting that $S(\operatorname{\mathcal{A}}_{q},\operatorname{\mathcal{P}},z)$ sifts out more elements of $\operatorname{\mathcal{A}}$ than $S(\operatorname{\mathcal{A}}_{q},\operatorname{\mathcal{P}},D_{q})$.
@@
$\leq k_{1}e^{-\gamma}\left(1+\frac{1}{2(\log\widetilde{D})^{2}}\right)\sum_{\begin{subarray}{c}q\in\operatorname{\mathcal{P}}\\ z\leq q<y\end{subarray}}\left(1-\frac{\log q}{\log y}\right)|\operatorname{\mathcal{A}}_{q}|\left(\frac{F(s_{q})+\varepsilon C_{1}(\varepsilon)e^{2}h(s_{q})}{{\log X}}\right)$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\sum_{\begin{subarray}{c}q\in\operatorname{\mathcal{P}}\\ z\leq q<y\end{subarray}}\left(1-\frac{\log q}{\log y}\right)\sum_{\begin{subarray}{c}d\mid P(z)\\ d<QD_{q}\end{subarray}}|r_{q}(d)|.$ (3.37)
@@
since $\operatorname{\mathcal{A}}$ is a set of consecutive integers. Now, because $k_{1}\leq 6$, we have $0<s_{q}\leq 3$ and thus by Lemma 3.2,
@@
$\qquad\leq\frac{1}{(2-3\alpha)\log X}\left(3\log\frac{k_{1}}{k_{2}}+(k_{2}(2-3\alpha)-3)\log\frac{2-3\alpha-\frac{3}{k_{2}}}{2-3\alpha-\frac{3}{k_{1}}}\right)$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\frac{5k_{1}^{4}}{k_{\alpha}(\log X)^{4}}\left(1-\frac{k_{2}}{k_{1}}\right).$
@@
$\sum_{\begin{subarray}{c}q\in\operatorname{\mathcal{P}}\\ z\leq q<y\end{subarray}}\left(1-\frac{\log q}{\log y}\right)|\operatorname{\mathcal{A}}_{q}|\left(\frac{F(s_{q})+\varepsilon C_{1}(\varepsilon)e^{2}h(s_{q})}{\log X}\right)\leq M_{1}(X)+M_{2}(X),$
@@
In this section we combine the results of Section 3 to prove Theorem 1.4. Here, the overarching approach is to use the computation in Theorem 1.3 for small $n$, and Lemma 3.6 for large $n$. From here onwards, the set $\operatorname{\mathcal{A}}$ is as defined in (3.1), $X$ is as in (3.12), and in the notation of Section 3, we set $k=2$, $k_{1}=6$, and $k_{2}=2.16$.
@@
For $a\geq 12$, the lemma follows directly from [12, Lemma 7]. The case $1<a\leq 12$} can then be verified via a short computation, noting that
@@
Our explicit version of ($Q_{0}(c,\delta)$) is then as follows. In particular, for $N>10^{40}$ we may take $c=0.2$ and $\delta=\frac{1}{4}$. This result could certainly be optimised further, however any improvement would have a negligible impact in our subsequent proof of Theorem 1.4.
@@
Let $N\geq 10^{40}$, $z=X^{1/6}$ and $y=X^{1/2.16}$. Then,
$\sum_{\begin{subarray}{c}z\leq q<y\\ q\in\mathcal{P}\end{subarray}}|\operatorname{\mathcal{A}}_{q^{2}}|\leq 0.2|\operatorname{\mathcal{A}}|^{3/4}.$
@@
$z\geq N^{1/6}=\frac{1}{3^{1/4}}(3N^{2/3})^{1/4}\geq\frac{1}{3^{1/4}}|\operatorname{\mathcal{A}}|^{1/4}.$
@@
$\frac{2.22|\operatorname{\mathcal{A}}|}{z\log z}\leq 2.22\cdot 3^{1/4}\frac{|\operatorname{\mathcal{A}}|^{3/4}}{\log(4.64\cdot 10^{6})}\leq 0.191|\operatorname{\mathcal{A}}|^{3/4}.$ (4.2)
@@
$\pi(y)\leq 1.04\frac{y}{\log y}\leq 0.025y\leq 0.025|\operatorname{\mathcal{A}}|^{0.69}.$ (4.3)
@@
$\sum_{\begin{subarray}{c}z\leq q<y\\ q\in\mathcal{P}\end{subarray}}|\operatorname{\mathcal{A}}_{q^{2}}|\leq 0.191|\operatorname{\mathcal{A}}|^{3/4}+0.025|\operatorname{\mathcal{A}}|^{0.69}\leq 0.2|\operatorname{\mathcal{A}}|^{3/4}$
@@
Let $N=n^{3}\geq 10^{40}$ throughout since smaller values of $N$ are covered by Theorem 1.3. We apply Lemma 4.3 and Lemma 3.6 with $k=2$, $z=X^{1/6}$, $y=X^{1/2.16}$}, and
$\lambda=k+1-k_{2}=0.84$
@@
$r_{2}(\operatorname{\mathcal{A}})\geq\frac{0.84}{2}S(\operatorname{\mathcal{A}},\operatorname{\mathcal{P}},z)-\frac{1}{2}\sum_{z\leq q\leq y}\left(1-\frac{\log q}{\log y}\right)S(\operatorname{\mathcal{A}}_{q},\operatorname{\mathcal{P}},z)-0.084|\operatorname{\mathcal{A}}|^{3/4}.$ (4.4)
@@
$0.084|\operatorname{\mathcal{A}}|^{3/4}\leq 10^{-5}\thinspace\frac{N^{2/3}}{\log X}.$ (4.5)
Next we give a lower bound for the first term in (4.4). Since $z=X^{1/6}>10^{5}$, applying Lemma 3.1 with Lemmas 3.3 and 3.4 gives
$S(\operatorname{\mathcal{A}},\operatorname{\mathcal{P}},z)$ $>|\operatorname{\mathcal{A}}|V(z)\cdot(f(s)-\varepsilon C_{2}(\varepsilon)e^{2}h(s))-\sum_{d<2D}\mu^{2}(d).$
@@
$f(s)-110\varepsilon e^{2}h(s)>C(s):=\frac{2e^{\gamma}\log(s-1)-0.093e^{2-s}}{s}.$ (4.7)
To bound the remainder term in (4), we apply the bound (see [19, Lemma 4.6])
@@
noting that $D=z^{s}\geq z^{3}\geq 10^{20}$. Substituting (4.7) and (4.8) into (4),
@@
Since $k_{1}=6$, we find that the lower bound in (4.9) is maximised around $s=3.54$ whereby $C(s)>0.93$ and
$\frac{0.84}{2}S(\operatorname{\mathcal{A}},\operatorname{\mathcal{P}},z)>3.9\frac{N^{2/3}}{\log X}.$ (4.10)
Finally we bound the second term in (4.4). In the notation of Proposition 3.9 we set $\alpha=0.06$ so that
$D_{y}=X^{\frac{2}{3}-0.06-\frac{1}{2.16}}>5\cdot 10^{5}$ (4.11)
and it is valid, via Lemma 3.4, to set $Q=2$ and $\varepsilon=2.8\cdot 10^{-4}$. In particular, Proposition 3.9 gives
$\frac{1}{2}\sum_{\begin{subarray}{c}q\in\operatorname{\mathcal{P}}\\ z\leq q<y\end{subarray}}\left(1-\frac{\log q}{\log y}\right)S(\operatorname{\mathcal{A}}_{q},\operatorname{\mathcal{P}},z)\leq\frac{6e^{-\gamma}}{2}\left(1+\frac{1}{2(\log D_{y})^{2}}\right)(M_{1}(X)+M_{2}(X))+\frac{\mathcal{E}(X)}{2}.$
@@
$3e^{-\gamma}\left(1+\frac{1}{(\log D_{y})^{2}}\right)\leq 1.7,$
$M_{1}(X)\leq\frac{2.04N^{2/3}}{\log X},$
$M_{2}(X)\leq(1.9\cdot 10^{-8})\frac{N^{2/3}}{\log X}$
@@
$r_{2}(\operatorname{\mathcal{A}})>0.27\frac{N^{2/3}}{\log X},$